OpenStudy (anonymous):

complete the ordered pairs so that each is a solution of the given linear equation. Then graph. x=3-y

OpenStudy (anonymous):



OpenStudy (anonymous):

x=3-y x-3=-y -x+3=y

OpenStudy (anonymous):

I have to complete the table in order to graph

OpenStudy (anonymous):

so how would I find out what y is in the table in order to find the points to the equation

OpenStudy (anonymous):

you plug in value for x and get the y

OpenStudy (anonymous):

okay but the equation is x=3-y that is why I am confused

OpenStudy (anonymous):

so the first one would be 0=3-y?

OpenStudy (anonymous):

yeah, solve for y

OpenStudy (anonymous):

so the first point would be y=-3?

OpenStudy (anonymous):

y==3

OpenStudy (anonymous):

 ?